OpenStudy (anonymous):

5x^3 + x^2 - 20x - 4

OpenStudy (anonymous):

factorize or solve = 0?

OpenStudy (anonymous):

factor

OpenStudy (anonymous):

ok, with x^3 u usually have to guess one of the solutions and then make polinomial division

OpenStudy (anonymous):

but thre might be another way for this one let me check 1 sec

OpenStudy (anonymous):



OpenStudy (anonymous):

(x-2)(5x^2+11x+2)

OpenStudy (anonymous):

(x-2)(5x^2+11x+2)

OpenStudy (anonymous):

(x-2) (x+2) (5 x+1)

OpenStudy (anonymous):

yes, is a double fact first asociate 5x³+x² and - (20x-4) so x²(5x+1)-4(5x+1)

myininaya (myininaya):

\[5x^3+x^2-20x-4=(5x^3+x^2)+(-20x-4) \] \[=x^2(5x+1)-4(5x+1)=(5x+1)(x^2-4)=(5x+1)(x-2)(x+2)\]